# Identification of Dynamic Nonlinear Panel Models

### v2

@@ 2.2 Identification Strategy
$=\Pr(Y_{it}=1,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st})+\Pr(z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}<a\mid z_{st})$
$=1-\Pr(Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st}).$
Given the lower and upper bound for the distribution $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$, we are ready to establish identifying conditions for $\theta_{0}$. Since Assumption 1 requires that $v_{t}\mid Z_{ist}\stackrel{{\scriptstyle d}}{{\sim}}v_{s}\mid Z_{ist}$, it implies the upper bound for the distribution of $v_{s}\mid Z_{ist}$ must be larger than the lower bound for the distribution of $v_{t}\mid Z_{ist}$ otherwise Assumption 1 cannot hold. This restriction characterizes an identified set for $\theta_{0}$, presented in the following proposition.
Proposition 1.
Under Assumption 1, an identified set $\Theta_{I,1}$ for $\theta_{0}$ is the set of parameters that satisfy the following conditions:
$1-\Pr(Y_{is}=0,z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma\geq a\mid z_{st})\geq\Pr(Y_{it}=1,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\leq a\mid z_{st}),$ (7)
for any $a\in\mathcal{R}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Proposition 1 characterizes an identified set for $\theta_{0}$, using the conditional joint distribution of $(Y_{it},X_{it})$ given $Z_{ist}$ across any pair of periods $(s,t)$. The identification result allows for endogeneity in the covariate $X_{ist}$ and does not rely on any excluded instruments. [^7] When a sub-vector $Z_{it}^{exc}$ of $Z_{it}$ is assumed to be excluded from the binary choice model, then the identifying restriction in (7) is adjusted by setting the coefficient of $Z_{it}^{exc}$ as zero: $\beta_{0}^{exc}=0$.
The identifying condition in (7) can be reformulated as conditional moment inequalities, by substituting the probability $\Pr(Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st})$ with the conditional expectation $\mathbb{E}\left[\mathbf{\mathbbm{1}}\{Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\}\mid z_{st}\right]$. Consequently, the estimation and inference for $\theta_{0}$ can be conducted using established methods such as Chernozhukov, Hong, and Tamer (2007), Andrews and Shi (2013), and Chernozhukov, Lee, and Rosen (2013).
@@
Our “partial stationarity” condition accommodates the “full stationarity” condition (2) as a special case. Under full stationarity, or in other words, when all covariates are exogenous, Proposition 1 specializes to the following “maximum-score-type” identifying restrictions in Manski (1987):
@@
Remark 2.
In the special case where the only endogenous covariate is one lagged outcome variable, as studied in Khan, Ponomareva, and Tamer (2023), the identifying restrictions yielded by Proposition 1 become
$1-\Pr(Y_{is}=0,z_{s}^{\prime}\beta_{0}+Y_{i,s-1}\gamma_{0}\geq a\mid z_{st})\geq\Pr(Y_{it}=1,z_{t}^{\prime}\beta_{0}+Y_{i,t-1}\gamma_{0}\leq a\mid z_{st}),$
for any $a\in\{z_{t}^{\prime}\beta_{0},z_{t}^{\prime}\beta_{0}+\gamma_{0},z_{s}^{\prime}\beta_{0},z_{s}^{\prime}\beta_{0}+\gamma_{0}\}$. As shown in Appendix A.4, the above identifying condition on $\theta_{0}$ implies the identifying conditions in Khan, Ponomareva, and Tamer (2023, Theorem 1, p. 8-9), under different values of $(z_{s}^{\prime}\beta_{0},z_{t}^{\prime}\beta_{0},\gamma_{0})$.
Proposition 1 does not impose restrictions on the support of the endogenous covariate $X_{it}$, allowing it to be either discrete or continuous. When $X_{it}$ is continuous, then Proposition 1 requires condition (7) to hold for any $a\in\mathcal{R}$ so it exploits the whole distribution information of $(X_{it},Y_{it})$ at any value $a$. For the estimation and inference of $\theta_{0}$, we can follow the conventional inference method to discretize the space of $\mathcal{R}$ and pick $K_{n}$ points for a. In the case of discrete $X_{it}$, the dimension of $a$ can be further reduced, and the identified set $\Theta_{I,1}$ can be characterized by finite number of restrictions, shown in the following corollary.
Corollary 1.
When the endogenous covariate $X_{it}\in\{a_{1},...,a_{K}\}$ only takes finite number of values, then the identified set $\Theta_{I,1}$ is the set of parameters $\theta$ that satisfy condition (7) for any $a\in\{z_{s}^{\prime}\beta+a_{k}^{\prime}\gamma,z_{t}^{\prime}\beta+a_{k}^{\prime}\gamma\}_{k=1}^{K}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Corollary 1 reduces the number of restrictions in Proposition 1 to $2K$ number of moment restrictions for any pair of two periods $s,t$. For the dynamic model where $X_{it}$ includes one lagged dependent variable $X_{it}=Y_{i,t-1}\in\{0,1\}$, there are only four conditional restrictions for $\theta_{0}$. When $X_{it}$ includes $j$ lagged dependent variable $X_{it}=(Y_{i,t-1},Y_{i,t-2},...Y_{i,t-j})$, then there will be $2^{j+1}$ restrictions.
We show that $\Theta_{I,1}$ is an identified set for $\theta_{0}$, however, it is still unclear whether condition (7) exploits all the available information for $\theta_{0}$. The following theorem establishes the sharpness of the identified set $\Theta_{I,1}$ when $X_{it}$ takes finite number of values.
Theorem 1 (Sharpness).
Under Assumption 1, the identified set $\Theta_{I,1}$ is sharp when $X_{it}$ takes finite number of values for any $t\leq T$.
Proposition 1 provides identification results without restrictions on the support of $X_{it}$, and Theorem 1 shows that the results already exhausted all available information for $\theta_{0}$ when $X_{it}$ only takes finite number of values. This results implies that for dynamic models with any number of lagged dependent variables, the identified set $\Theta_{I,1}$ is the smallest set we can obtain.

### v3

@@ 2.2 Key Identification Strategy
which is equivalent to
$\mathbf{\mathbbm{1}}\left\{v_{it}\leq c\right\}\leq 1-\left(1-Y_{it}\right)\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c\right\}.$ (8)
Taking conditional expectations of (7) and (8) given $Z_{i}=z\equiv\left(z_{t}\right)_{t=1}^{T}$,
$\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c\right|Z_{i}=z\right)$
[…]
=\ = $\mathbb{P}\left(\left.v_{is}\leq c\right|Z_{i}=z\right)$
[…]
where the middle equality follows from the partial stationarity condition (Assumption 1). Specifically, observe that Assumption 1 implies the partial stationarity of $v_{it}$ given $Z_{i}$, i.e.,
$v_{it}\mid Z_{i}\stackrel{{\scriptstyle d}}{{\sim}}v_{is}\mid Z_{i},$
so that $\mathbb{P}\left(\left.v_{it}\leq c\right|Z_{i}=z\right)=\mathbb{P}\left(\left.v_{is}\leq c\right|Z_{i}=z\right)$ for any $c\in\mathcal{R}$.
Essentially, in the above we exploit the joint occurrence of $v_{it}\leq Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}$ and $Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c$ to deduce an implication on the composite error $v_{it}\leq c$ that is free of the endogenous covariates $X_{it}$, and then leverage the partial stationarity of $v_{it}$ given $Z_{i}$ for intertemporal comparisons.
Since the lower and upper bounds in (9) hold for any $t$ and $s$, we summarize the identifying restrictions (9) across all time periods in the following proposition.
Proposition 1 (Identified Set).
Under model (5) and Assumption 1, $\theta_{0}\in\Theta_{I}$, where the identified set $\Theta_{I}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
$\max_{t=1,...,T}\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right|Z_{i}=z\right)\leq 1-\max_{s=1,...,T}\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq c\right|Z_{i}=z\right)$ (10)
for any $c\in\mathcal{R}$ and any realization $z=\left(z_{1},...,z_{T}\right)$ in the support of $Z_{i}$.
Proposition 1 characterizes the identified set $\Theta_{I}$ for $\theta_{0}$ as restrictions on the conditional joint distribution of $Y_{it}$ and $X_{it}$ given $Z_{i}=z$. More specifically, the restrictions in (10) can be regarded as a collection of conditional moment inequalities that relate $\mathbf{\mathbbm{1}}\left\{Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right\}$ and $\mathbf{\mathbbm{1}}\left\{Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right\}$ conditional on $Z_{i}=z$.
Proposition 1 holds regardless of whether the endogenous covariates $X_{it}$ are discrete or continuous. When $X_{it}$ are continuous (taking a continuum of values), then Proposition 1 requires that condition (10) hold for a continuum of constants $c\in\mathcal{R}$, so that (the information in) the whole joint distribution of the binary variable $Y_{it}$ and the continuous variable $z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma$ can be captured by the collection of joint distributions of $\left(Y_{it},\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right\}\right)$ across all possible cutoff values $c$.
However, when $X_{it}$ are discrete, such as in the AR$\left(p\right)$ dynamic model where $X_{it}$ consists of $p$ lagged binary outcome variables, there is no need to evaluate (10) at all possible values of $c\in\mathcal{R}$, since the inequalities in (10) can only bind at finitely many values of $c$. We formalize this observation via the following Proposition.
Proposition 2 (Identified Set with Discrete Endogenous Covariates).
Suppose that the endogenous covariate $X_{it}$ can only take finite number of values in $\left\{\overline{x}_{1},...,\overline{x}_{K}\right\}$ across all time periods $t=1,...,T$, then $\Theta_{I}=\Theta_{I}^{disc},$ where $\Theta_{I}^{disc}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ that satisfy condition (10) for any
$c\in\left\{z_{t}^{{}^{\prime}}\beta+\overline{x}_{k}^{{}^{\prime}}\gamma:k=1,...,K,t=1,...,T\right\},$ (11)
for any $z=\left(z_{1},...,z_{T}\right)$ in the support of $Z_{i}$.
Proposition 2 shows that the discreteness of the endogenous covariates $X_{it}$ help reduce the infinite number of inequality restrictions in Proposition 1 to finitely many, or more precisely, $KT$ ones (conditional on $z$).
The case of discrete $X_{it}$ is conceptually important, since it nests the dynamic AR$\left(p\right)$ model widely studied in the literature as a special case. Clearly, when $X_{it}$ consists of $p$ (finitely many) lagged binary outcome variables $Y_{i,t-1},...,Y_{i,t-p}$, then $X_{it}$ by construction can only take $K=2^{p}$ discrete values. Specialized further to the AR(1) model in KPT, Proposition 2 shows that the identified set $\Theta_{I}$ is characterized by $2T$ conditional restrictions, which is drastically smaller than the $9T\left(T-1\right)$ conditional restrictions listed out in KPT (even when $T$ is small).
@@
Following up on Remark 4, if pairwise partial stationarity is adopted, then Propositions 1 and 2 continue to hold with (10) adapted to the following “pairwise” version:
$\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right|Z_{i,ts}=z_{ts}\right)\leq 1-\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq c\right|Z_{i,ts}=z_{ts}\right),\forall\,(t,s),$ (12)
where $Z_{i,ts}:=(Z_{it},Z_{is})$ and $z_{ts}:=(z_{t},z_{s}).$ Relative to (10), the statement in (12) reflects the fact that pairwise partial stationarity is imposed on all pairs of time periods separately instead of all $T$ time periods jointly. It is straightforward to verify that the identification arguments above, in particular (7)-(9), carry over with all conditional probabilities/expectations taken conditional on $Z_{i,ts}=z_{ts}$ instead of $Z_{i}=z$.
2.3 Sharpness of Identified Set
So far we have only shown that $\Theta_{I}$, which equals to $\Theta_{I}^{disc}$ under discreteness of $X_{it}$, is a valid identified set for $\theta_{0}$. However, it is not yet clear whether it has incorporated all the available information for $\theta_{0}$. The following theorem establishes the sharpness of the identified set $\Theta_{I}$ under the discreteness of $X_{it}$.
Theorem 1 (Sharpness under Discreteness).
When $X_{it}\in\left\{\overline{x}_{1},...,\overline{x}_{K}\right\}$ for all $t=1,...,T$, then under model (5) and Assumption 1, the identified set $\Theta_{I}$ is sharp.
The formal definition of sharpness, along with the complete proof of Theorem 1, are available in Appendix A.2. In short, we show (by construction) that, for each $\theta\in\Theta_{I}\backslash\left\{\theta_{0}\right\}$, there exists a data generating process (DGP) that satisfies Assumption 1 and produces the same joint distribution of observable data $\left(Y_{i},X_{i},Z_{i}\right)$ under model (5) and the discreteness of $X_{it}$. Theorem 1 demonstrates that our key identification strategy based on the bounding of (endogenous) parametric index by arbitrary constants, as described in Section 2.2, is able to extract all the available information for $\theta_{0}$ from the model and the observable data, and thus it is impossible to further differentiate $\theta_{0}$ from alternatives in the identified set $\Theta_{I}$ under model (5) and our assumption of partial stationarity (without further restrictions).
Theorem 1 immediately implies that, in the special case of dynamic AR$\left(p\right)$ models where $X_{it}$ consists of discrete lagged outcomes, our characterization of the identified set $\Theta_{I}^{disc}$ in Proposition 2 is sharp. In particular, our result generalizes the corresponding result in KPT, which focuses on the AR(1) model. Furthermore, KPT characterizes the sharp identified set via $9T\left(T-1\right)$ conditional restrictions, the derivation of which is based on an exhaustive enumeration of lagged outcome realizations $Y_{i,t-1}$. In this paper we adopt an entirely different (and much more general) identification strategy, and arrive at a characterization of the identified set by $2T$ conditional restrictions, which we also show to be sharp by Theorem 1. Since our model and assumption specialize exactly to that in KPT under the AR(1) specification, it follows that our $2T$ restrictions must be able to reproduce all the $9T\left(T-1\right)$ restrictions in KPT. This demonstrates that our identification strategy not only applies more generally than the one in KPT, but also leads to a more elegant characterization of the sharp identified set with much fewer restrictions. We provide a more detailed explanation about this point in the next subsection.
Another conceptually remarkable, or surprising, feature of Proposition 2 and Theorem 1 is that they are established without reference to the exact nature, or interpretation, of the endogenous covariates $X_{it}$. The identified set $\Theta_{I}$ we characterized is valid and sharp regardless of whether $X_{it}$ are specified as lagged outcome variables or contemporarily endogenous covariates, as long as $X_{it}$ is discrete.
Our proof of sharpness consists of two main lemmas. First, we show for each $\theta\in\Theta_{I}\backslash\left\{\theta_{0}\right\}$ how to construct the per-period marginal distributions of errors that match the per-period marginal choice probabilities. Second, we show how to combine the $T$ per-period marginal distributions into an all-period joint distribution that matches the all-period joint choice probabilities, so that observational equivalence holds.
The proof techniques we exploited are also different from, and thus novel relative to, those used in the related work that leverages stationarity-type conditions for partial identification, such as Pakes and Porter (2022) for static multinomial choice model and KPT for dynamic AR(1) model. Instead of showing existence only, we provide a more explicit construction of the joint distribution of the latent variables, which is valid regardless of the exact type of endogeneity in $X_{it}$. In particular, a key challenge in proving sharpness based on stationarity-type conditions lies in that stationarity imposes only aggregate restrictions (via integrals/sums) of the joint distribution of errors, which is rather implicit to work with. A key innovation in our proof technique is to show how marginal/aggregate stationarity restrictions and joint choice probability restrictions can be satisfied simultaneously by an explicit, simple, and general construction, which might be of independent and wider use.
2.4 Reconciliation with Related Work
Our identifying restrictions in (10) and (11) have a somewhat “nonstandard” representation in terms of (conditional) joint probabilities instead of $Y_{it}$ and $X_{it}$ (given $Z_{i}$), instead of conditional probabilities of $Y_{it}$ given $X_{it}$ (such as lagged outcomes), which are more usually found in the related literature. Hence, we provide a more detailed discussion about the content and interpretation of our identifying restrictions, as well as a more explicit explanation of how they relate to exisiting results in the related literature.
Reconciliation with Manski (1987)
Consider first the special case where there are no endogenous covariates $X_{it}$, or in other words, $X_{it}$ is degenerate. In this case, our “partial stationarity” condition specializes to the “full stationarity” condition (2) as in Manski (1987). However, our identifying restriction (10) still has a different form than the identifying restriction in Manski (1987). To illustrate, focus on any two periods $\left(t,s\right)$, and observe that in this case our identifying restriction takes the form of
$\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta_{0}\leq c\right|z\right)\leq 1-\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta_{0}\geq c\right|z\right),\ \forall c,$ (13)
while the “maximum-score-type” identifying restrictions in Manski (1987) are of the form
$z_{s}^{{}^{\prime}}\beta_{0}\geq z_{t}^{{}^{\prime}}\beta_{0}\ \Leftrightarrow\ \mathbb{P}\left(\left.Y_{is}=1\right|z\right)\geq\mathbb{P}\left(\left.Y_{it}=1\right|z\right).$ (14)
The “maximum-score-type” identifying restriction (14) has a quite intuitive and interpretable representation: across two periods $\left(t,s\right)$ under full stationarity, the conditional choice probability at period $s$ is larger if and only if the index $z_{s}^{{}^{\prime}}\beta_{0}$ is larger. In contrast, our restriction (13) has a somewhat twisted representation even in this simple setting.
However, a closer look reveals that our (13) is exactly equivalent to Manski’s “maximum-score-type” identifying restrictions in the current context. To see this, notice that, by setting $c=z_{t}^{{}^{\prime}}\beta_{0}$ in (13), we obtain
$\mathbb{P}\left(\left.Y_{it}=1\right|z\right)=\,$ $\mathbb{P}\left(\left.Y_{it}=1\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}\leq z_{t}^{{}^{\prime}}\beta_{0}\right\}$
@@ 2.4 Reconciliation with Related Work
Hence, if $z_{s}^{{}^{\prime}}\beta_{0}\geq z_{t}^{{}^{\prime}}\beta_{0}$, i.e., the left-hand side of (14) holds, then the above implies that
$\mathbb{P}\left(\left.Y_{it}=1\right|z\right)\leq 1-\mathbb{P}\left(\left.Y_{is}=0\right|z\right)=\mathbb{P}\left(\left.Y_{is}=1\right|z\right),$
which becomes exactly the right-hand side of (14). Switching $t$ with $s$ in the argument above produces the other implication $z_{s}^{{}^{\prime}}\beta_{0}\leq z_{t}^{{}^{\prime}}\beta_{0}$Rightarrow $\mathbb{P}\left(\left.Y_{is}=1\right|z\right)\leq\mathbb{P}\left(\left.Y_{it}=1\right|z\right)$. Together these exactly constitute the “if-and-only-if” restriction in (14). Hence, even though our inequality restriction (13) looks different from the more intuitive “maximum-score-type” restriction, they both incorporate the same information.
Reconciliation with KPT (Khan, Ponomareva, and Tamer, 2023)
Now, consider the dynamic AR(1) model as studied in KPT, where the only endogenous covariate is the one-period lagged outcome variable, i.e., $X_{it}:=Y_{i,t-1}$.
To illustrate, first focus on any two periods $\left(t,s\right)$ only, and observe that in this case our identifying restriction becomes
$\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta_{0}+Y_{i,t-1}\gamma_{0}\leq c\right|z\right)\leq 1-\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta_{0}+Y_{i,t-1}\gamma_{0}\geq c\right|z\right),\ \forall c.$ (15)
Under the same model and assumption, KPT derives the following 9 inequality implications for $\left(t,s\right)$: [^6] We adapt the notation in KPT to our current notation, and state these 9 inequalities as strict inequalities, which lead to a simpler and more focused explanation. The equivalence between our restriction and the KPT restrictions still hold if their inequalities are stated in the weak form.
KPT(i): $\mathbb{P}\left(\left.Y_{it}=1\right|z\right)>\mathbb{P}\left(\left.Y_{is}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}+\left|\gamma_{0}\right|>0.$
KPT(ii): $\mathbb{P}\left(\left.Y_{it}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{i,s}=0,Y_{i,s-1}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}-\min\left\{0,\gamma_{0}\right\}>0.$
KPT(iii): $\mathbb{P}\left(\left.Y_{it}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{i,s}=0,Y_{i,s-1}=0\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}+\max\left\{0,\gamma_{0}\right\}>0.$
KPT(iv): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=1\right|z\right)>\mathbb{P}\left(\left.Y_{is}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}+\max\left\{0,\gamma_{0}\right\}>0.$
KPT(v): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0,Y_{i,s-1}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}>0.$
KPT(vi): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0,Y_{i,s-1}=0\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}+\gamma_{0}>0.$
KPT(vii): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=0\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}-\min\left\{0,\gamma_{0}\right\}>0.$
KPT(viii): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=0\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0\text{$,$}Y_{i,s-1}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}-\gamma_{0}>0.$
KPT(ix): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=0\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0,Y_{i,s-1}=0\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}>0.$
In a way, the 9 inequality restrictions in KPT above are similar to the “maximum-score restrictions”, in the sense that all of them take the form of logical implications between intertemporal comparisons of various conditional probabilities and intertemporal differences of covariate indexes.
Using a very different identification strategy than the one in KPT, we arrived at our inequality restriction (15), which looks very different from the collection of 9 inequality restrictions in KPT. At first sight it is not clear how (15) relates to and compares with the 9 KPT restrictions. However, a closer look again reveals that our restriction (15) can reproduce all the 9 restrictions in KPT, and thus incorporate all the information therein in a unified format.
Take KPT(v) as an illustration and suppose that the left-hand side of KPT(v) holds, then it implies
$\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0,Y_{i,s-1}=1\right|z\right).$ (16)
With $X_{it}=Y_{i,t-1}$, our inequality restriction (15) can be equivalently rewritten as follows,
$\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=1\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0}\leq c\right\}+\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=0\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}\leq c\right\}$
[…]
where the realization of $Y_{i,t-1}$ is explicitly enumerated as in KPT.
Note that we can further relax condition (17) by dropping the two probabilities $\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=0\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}\leq c\right\}$ and $\mathbb{P}\left(\left.Y_{is}=0,\ Y_{i,s-1}=0\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{s}^{{}^{\prime}}\beta_{0}\geq c\right\}$ as it makes the lower bound smaller and the upper bound larger:
$\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=1\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0}\leq c\right\}\leq\ \\ 1-\mathbb{P}\left(\left.Y_{is}=0,\ Y_{i,s-1}=1\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{s}^{{}^{\prime}}\beta_{0}+\gamma_{0}\geq c\right\}.$
Then, the statement that $\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0}\leq c\right\}$ and $\mathbf{\mathbbm{1}}\left\{z_{s}^{{}^{\prime}}\beta_{0}+\gamma_{0}\geq c\right\}$ both holds is precisely equivalent to the following statement of
[…]
By contraposition, it leads to exactly the same implication of KPT(v):
[…]
Hence, we have shown that (17) implies KPT(v).
Similarly, it is shown in Appendix A.3 that (17) implies all 9 restrictions in KPT. In fact, the representation (17) reveals why there are precisely 9 KPT-type restrictions. The two period-$t$ indicators $\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0}\leq c\right\}$ and $\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}\leq c\right\}$ in the upper expression of (17) may take 3 “useful” [^7] The 4th combination, $\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0}\leq c\right\}=\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0}\leq c\right\}=0$, will make the upper expression of (17) equal to 00, so that the inequality (17) holds trivially. Hence, this $\left(0,0\right)$ combination is not useful. combinations $\left(1,0\right),\left(0,1\right)$ and $\left(1,1\right)$, while the two period-$s$ indicators $\mathbf{\mathbbm{1}}\left\{z_{s}^{{}^{\prime}}\beta_{0}+\gamma_{0}\geq c\right\}$ and $\mathbf{\mathbbm{1}}\left\{z_{s}^{{}^{\prime}}\beta_{0}\geq c\right\}$ in the lower expression of (17) may also take 3 useful combinations. Consequently, in total there are $3\times 3=9$ useful combinations, which exactly correspond to the $9$ left-hand-side suppositions in the 9 KPT restrictions.
Hence, while our restriction (15) appears very different from the 9 KPT restrictions, it actually automatically incorporates all the KPT restrictions. In particular, by treating the endogenous covariate $X_{it}=Y_{i,t-1}$ as a random variable, our restriction (15) automatically aggregates the identifying information across all possible realizations of $Y_{i,t-1}$, without the need to explicitly consider each possibility separately.
Now, consider a general setting with $T\geq 2$ periods. By our Proposition 2 and Theorem 1, the sharp identified set can be characterized by $2T$ restrictions, which are generated by evaluating (10) at each $c$ of the $2T$ points in $\left\{z_{t}^{{}^{\prime}}\beta,z_{t}^{{}^{\prime}}\beta+\gamma:t=1,...,T\right\}$. In contrast, across $T$ periods the KPT approach produces $9T\left(T-1\right)$ restrictions, which are generated by imposing the $9$ KPT restrictions across all ordered time pairs $\left(t,s\right)$. Hence, our approach provides a much simpler characterization of the sharp identified set, using a significantly smaller number of restrictions. For example, with $T=2$ periods, we have $4$ restrictions while KPT has $18$; with $T=3$, we have $6$ restrictions while KPT has 54. Hence, the reduction in the number of restrictions relative to KPT is quite remarkable.
In summary, while the representation of our identifying restrictions in (10) and (11) may appear somewhat unnatural in the first place, it actually becomes equivalent to more familiar (and intuitive) representations in the specialized settings of Manski (1987) and KPT.